# Time-Dependent Black Hole Lensing and the Unified Weak-to-Strong Deflection Framework

### v1

@@ IV.2 Time delays, multiplicity, and photon-ring modulation
$=T_{\mathrm{whirl}}(t_{o})+\frac{2\pi}{\bar{a}(t_{o})}\,\tilde{a}(t_{o})+\ldots,$ (48)
where the second term arises from the logarithmic piece through $\ln s_{n+1}-\ln s_{n}=-2\pi/\bar{a}(t_{o})$. Linearizing Eq. (48) using Eqs. (39) and (47) shows that $\Delta T_{n+1,n}$ acquires a harmonic modulation $\propto\mathrm{Re}\,{e^{-i\omega t_{o}}}$ with an amplitude set by $\gamma_{1},\tilde{a}_{1},a_{1}$.
Figure 6: Inter-image time delays $\Delta T_{n+1,n}(t_{o})$ from Eq. (46). The baseline separation is set by the whirl period $T_{\rm whirl}=2\pi/\Omega_{\rm ph}$ with $\Omega_{\rm ph}=(3\sqrt{3}\,M)^{-1}$, while the weak modulation arises from the time-dependent coefficients in Eq. (47) and the $\bar{a}(t_{o})$ dependence. The delays oscillate in phase with the spacing modulation $q(t_{o})$, reflecting their common origin in the near-photon-sphere kernel.
Figure 6 makes the phase-locked behavior explicit: variations in $\Delta T_{n+1,n}(t_{o})$ track the same QNM phase that controls $q(t_{o})$ and $\theta_{\rm ring}(t_{o})$. Eliminating $t_{o}$ between Eqs. (48) and (50) yields the differential correlation summarized in Eq. (62), further underscoring the unified kernel control.
Finally, the photon ring corresponds to the accumulation of the sequence $\theta_{n}(t_{o})$ as $n\to\infty$ Gralla et al. (2019). From Eq. (43) we obtain the instantaneous ring radius
@@
$=\theta_{c}\left[1+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\beta_{1}\right\}\right],$ (49)
@@
$=e^{-2\pi}\left[1+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{2\pi\,a_{1}}{\bar{a}^{2}}\right\}\right]+\mathcal{O}(\varepsilon^{2}),$ (50)
where we used $\bar{a}=1$ for Schwarzschild in the static limit. Thus, both the ring radius and the logarithmic spacing inherit the QNM frequency and phase from Eq. (24). The same analysis yields the magnifications $\mu_{n}(t_{o})$, which scale as $\mu_{n}\propto q(t_{o})^{n}$ up to a slowly varying prefactor; we refrain from writing the prefactor explicitly because it depends on the source configuration, while the time-dependent exponential factor follows directly from Eq. (50). Related visibility–domain signatures and detectability of higher-order rings in realistic accretion models are discussed in Vincent et al. (2022); an analytic “gap parameter” characterization of higher-order rings in spherically symmetric metrics is given in Aratore et al. (2024).
@@
Eqs. (43)-(50), together with the deflection law in Eq. (38), provide a self-contained, analytic description of how a single even-parity $\ell=2,m=0$ QNM coherently imprints on the arrival times, multiplicity, and morphology of strong-field images, culminating in a phase-locked modulation of the photon ring.
@@ V Imaging Diagnostics and Continuity Across Regimes
We now assemble a single, observer-based description of lensing observables that is valid from the weak field to the time-dependent strong-deflection limit. The central object is the Born line-of-sight map in Eq. (24), which yielded the far-field expansion of Eq. (30) and the strong-field law of Eq. (38). Our task is twofold: first, to show that these limits admit a common overlap where they agree term by term once expressed on the screen defined in Sec. II.2; second, to translate the matched deflection into gauge-safe imaging diagnostics for centroid motion, multi-image structure, and photon-ring modulations.
@@ V.1 The unified matching
The matching proceeds directly at the level of the master integral in Eq. (24). We split the integral at an arbitrary radius $r_{m}$ chosen so that $3M<r_{m}\ll r_{o}$, writing
@@
$+\int_{\lambda(r_{m})}^{\lambda_{o}}d\lambda\,\tilde{\mathcal{K}}_{A}\,e^{i\omega\tau}\equiv\mathcal{A}_{A}^{\text{in}}+\mathcal{A}_{A}^{\text{out}},$ (51)
@@
For $b\to b_{c}^{+}$ at fixed $r_{m}$, the inner integral reproduces the logarithmic structure derived in Eq. (37), hence
@@
$+\mathcal{D}_{A}(\omega;r_{m})\Bigr]\biggr\}+\ldots,$ (52)
with $\mathcal{C}_{A}$ independent of $r_{m}$ and $\mathcal{D}_{A}$ carrying the $r_{m}$-dependence required to cancel that of the outer piece. Conversely, for $M/b\ll 1$ at fixed $r_{m}$, the outer contribution matches the far-field form in Eq. (27),
$\delta\alpha_{A}^{\text{out}}(b,t_{o})=\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{c_{A}(\omega;r_{m})}{b}\,\mathcal{F}_{A}(\omega b)\right\}+\mathcal{O}\,\left(\frac{M}{b^{2}}\right),$ (53)
where $c_{A}(\omega;r_{m})$ depends on the cutoff but $\mathcal{F}_{A}$ does not. Adding Eqs. (52)-(53) yields a representation that is uniformly valid for $b>b_{c}$.
Figure 8: Kernel split and $r_{m}$ independence of the sum. The master line-of-sight amplitude is decomposed as in Eq. (51) into the inner (near-photon-sphere) piece and the outer (far-zone) piece. Each contribution depends on the arbitrary split radius $r_{m}$ and carries complementary logarithmic/finite parts, yet their sum is independent of $r_{m}$. This illustrates that physical observables derived from the Born kernel are unique and do not depend on the cutoff.
Figure 8 shows how the $r_{m}$-dependence cancels between the inner and outer integrals in Eq. (51). This cancellation underlies the robustness of the matched coefficients in Eq. (54) and, ultimately, the overlap relation in Eq. (55).
To make contact with the time-dependent SDL coefficients of Eq. (39), we expand the outer factor $\mathcal{F}_{A}(\omega b)$ for $b$ close to $b_{c}$. Using the standard near-critical relation between the longitudinal coordinate and the azimuthal angle, Eq. (36), one finds that $\mathcal{F}_{A}$ carries a logarithmic dependence identical to that of the inner piece. The $r_{m}$-dependence cancels between $\mathcal{D}_{A}(\omega;r_{m})$ and $c_{A}(\omega;r_{m})$, and the sum can be cast as
@@
$+\frac{\beta_{1,A}(\omega)}{\frac{b}{b_{c}}-1}+b_{1,A}(\omega)\Bigr]\biggr\}+\ldots,$ (54)
which is the vectorial version of Eq. (41) and fixes, component by component, the response coefficients $a_{1,A},\beta_{1,A},b_{1,A}$ in terms of the same kernel $\tilde{\mathcal{K}}_{A}$ used in the far-field calculation.
A practical matching condition follows by selecting an overlap domain where both asymptotics are valid. Let $b=b_{c}(1+\Delta)$ with $0<\Delta\ll 1$ and keep $M/b\ll 1$. Expanding Eq. (30) to first order in $\Delta$ and equating to Eq. (54) yields
@@
$+\bm{\Xi}(\omega),$ (55)
where $\bm{\mathcal{A}}=(c_{X}\mathcal{F}_{X},c_{Y}\mathcal{F}_{Y})$ from Eq. (30), $\hat{\bm{e}}_{b}$ is the screen radial unit vector, and $\bm{\Xi}(\omega)$ collects finite terms that are independent of the arbitrary split radius $r_{m}$. Expanding $A(b,\omega)$ about $b=b_{c}(1+\Delta)$ generates the $\ln\Delta$ and $1/\Delta$ structures that match the strong-deflection expansion on the inner side, making the regulator independence manifest. Eq. (55) states that the $1/b$ Born amplitude evaluated at $b_{c}$ encodes, through its singular $\Delta$-expansion, the three time-dependent SDL responses in Eq. (39). In particular, the residue of the $1/\Delta$ pole fixes $\beta_{1}(\omega)$ and therefore the photon-ring modulation in Eq. (49), while the coefficient of $\ln\Delta$ fixes $a_{1}(\omega)$ and hence the spacing modulation in Eq. (50).
Figure 9: Matching weak and strong limits in the overlap domain. The weak-field Born expression (Eq. (30), dashed) and the linearized SDL law (Eq. (41), solid) are compared at $b=b_{c}(1+\Delta)$ for small $\Delta>0$. Around an intermediate reference $\Delta$ the two curves coincide, and they agree within a band across the overlap window, illustrating Eq. (55): the near-critical expansion of the $1/b$ Born amplitude encodes the SDL coefficients, in particular the $1/\Delta$ residue that fixes $\beta_{1}(\omega)$.
Figure 9 visualizes the content of Eq. (55): the same kernel controls both limits, and their expansions meet smoothly for small but finite $\Delta$. The agreement pins down the triplet $(\beta_{1},a_{1},b_{1})$ that governs ring-radius and spacing modulations (Sec. (IV)).
Finally, since Eqs. (24), (30), and (38) are all constructed using the observer tetrad of Eq. (9), the matching preserves the gauge safety guaranteed by Eq. (16). The result is a single kernel-based description in which weak-field centroid wobble, intermediate-field image shifts, and strong-field ring modulations are different faces of the same frequency-resolved response.
@@ V.2 Screen-plane observables: centroid, multi-images, and ring
All observables are defined on the screen introduced in Eq. (13) and inherit their time dependence from the master kernel in Eq. (24). We collect here the gauge-safe quantities that diagnose the weak, intermediate, and strong lensing regimes and emphasize how they are continuously related by the matching established in Sec. V.1.
@@
$\zeta\equiv\frac{X+iY}{r_{o}},$ (56)
so that a ray of impact parameter $b$ arriving at $(X,Y)$ corresponds to $\zeta(b,t_{o})=\alpha_{X}(b,t_{o})/r_{o}+i\alpha_{Y}(b,t_{o})/r_{o}$, with $\bm{\alpha}$ given by Eq. (19). For a narrow bundle around $b$ the centroid angle is the intensity-weighted mean,
$\zeta_{c}(t_{o})=\zeta_{c,0}+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\mathcal{B}(\omega)\right\}+\mathcal{O}\,\left(\frac{M}{r_{o}}\frac{M}{b^{2}}\right),$ (57)
which is the complex form of Eq. (32). The complex amplitude $\mathcal{B}(\omega)$ depends on the bundle’s impact-parameter distribution through $\mathcal{A}_{A}(b,\omega)$ in Eq. (24). Since both the phase and modulus of $\mathcal{B}$ are determined by the same kernel that appears in the strong-field coefficients of Eq. (39), the centroid “wobble” is phase-locked to the modulations of the relativistic images and of the photon ring discussed below. In particular, defining
$\Phi_{\rm c\,-\,ring}\equiv\arg\mathcal{B}-\arg\beta_{1},$ (58)
the relative phase between the weak-field centroid wobble and the strong-field ring-radius oscillation, Eq. (49), is an overlap-domain diagnostic of the common QNM driver.
Figure 10: Phase-lock diagnostic across regimes. We show the phases associated with the weak-field centroid wobble (from Eq. (32)) and the strong-field ring-radius modulation (from Eq. (49)), together with their difference $\Phi_{\rm c\!-\!ring}$ defined in Eq. (58). The two phases advance in parallel at frequency $\omega_{R}$ while their difference remains constant, demonstrating that weak- and strong-field screen observables are driven by the same QNM kernel.
Figure 10 exhibits the anticipated phase locking: the centroid and ring phases are parallel lines with a constant offset $\Phi_{\rm c\!-\!ring}$. This provides a direct, gauge-safe confirmation that the Born kernel controls both observables, as argued in Sec. V.1 and encoded in Eq. (55).
Relativistic images are labeled by the winding number $n\in\mathbb{N}$ and obey Eq. (43). Their instantaneous angular radii $\theta_{n}(t_{o})$ separate naturally into a ring-centered offset,
@@
with $\theta_{\rm ring}(t_{o})$ given by Eq. (49). The complex offset amplitude,
$\delta\theta_{n}^{\rm off}=\theta_{c}\,s_{n}^{(0)}\,\left(\frac{b_{1}}{\bar{a}}-\frac{\bar{b}-2\pi n}{\bar{a}^{2}}\,a_{1}\right),$ (60)
is the second term in Eq. (44). The geometric decay $\propto s_{n}^{(0)}$ ensures that higher-order images accumulate at the ring while remaining coherently modulated at the QNM frequency $\omega$.
The logarithmic spacing is governed by $q(t_{o})$ in Eq. (50). Writing $q(t_{o})=q_{0},[1+\varepsilon\,\mathrm{Re}{e^{-i\omega t_{o}}\,\delta_{q}}]$ with $q_{0}=e^{-2\pi/\bar{a}}$ and $\bar{a}=1$ for Schwarzschild, we have
$\delta_{q}=\frac{2\pi\,a_{1}}{\bar{a}^{2}},$ (61)
so that a measurement of the instantaneous ratio $\left(\Delta\theta_{n+1}/\Delta\theta_{n}\right)(t_{o})$ yields $\mathrm{Re}{e^{-i\omega t_{o}}\delta_{q}}$ directly. In combination with Eq. (49), which determines $\mathrm{Re}{e^{-i\omega t_{o}}\beta_{1}}$, this provides two independent strong-field observables tied to the same kernel coefficients $a_{1}$ and $\beta_{1}$.
Arrival times follow from the travel-time function $T(b,t_{o})$ in Eq. (45). Evaluated at $b_{n}(t_{o})$ from Eq. (42), the inter-image delays $\Delta T_{n+1,n}(t_{o})$ are given by Eq. (48) and thus oscillate at the QNM frequency through the modulations in Eq. (47). Eliminating $t_{o}$ between Eqs. (48) and (50) gives a purely geometric correlation,
$\frac{d}{dt_{o}}\ln q(t_{o})=\frac{2\pi}{\bar{a}^{2}}\,\frac{d}{dt_{o}}\bar{a}(t_{o})=\mathcal{C}_{T}\,\frac{d}{dt_{o}}\Delta T_{n+1,n}(t_{o})+\mathcal{O}(\varepsilon^{2}),$ (62)
where $\mathcal{C}_{T}$ is a constant fixed by the static Schwarzschild values of $\tilde{a}$ and $\bar{a}$. Eq. (62) expresses the common origin of spacing and delay modulations in the near-photon-sphere kernel and, in our framework, is the time-domain analogue of the matching relation (55).
@@
$\mu_{n+1}(t_{o})\simeq q(t_{o})\,\mu_{n}(t_{o}),$ (63)
up to a slowly varying prefactor that is independent of $n$. Using Eq. (50), the fractional modulation of $\mu_{n}$ at fixed $n$ is therefore $\delta\mu_{n}/\mu_{n}\simeq\mathrm{Re}{e^{-i\omega t_{o}}\delta_{q}}$, while at fixed $t_{o}$ the ratio $\mu_{n+1}/\mu_{n}$ isolates $q(t_{o})$. Together with Eqs. (49)-(50), this yields a closed set of strong-field observables $\theta_{\rm ring},q,\mu_{n+1}/\mu_{n}$} that are all phase-locked to $\omega$.
Finally, the weak- and strong-field diagnostics can be combined without reference to the arbitrary matching radius used in Eq. (51). A convenient pair is
$\left(\zeta_{c}(t_{o})-\zeta_{c,0}\right)\quad\text{and}\quad\theta_{\rm ring}(t_{o})-\theta_{c},$ (64)
whose complex phases are $\arg\mathcal{B}$ and $\arg\beta_{1}$, respectively. The constancy of $\Phi_{\rm c\,-\,ring}$ in Eq. (58) across the overlap domain provides a direct test of continuity from Eq. (30) to Eq. (38). Because all quantities in Eqs. (57)-(64) are defined through the tetrad map Eq. (13), the gauge-independence argument of Eq. (16) applies verbatim: centroid motion, inter-image spacing, ring radius, and magnification ratios are bona fide $\mathcal{O}(\varepsilon)$ observables tied to a single, unified response encoded in the Born kernel (24).

### v2

@@ IV.2 Time delays, multiplicity, and photon-ring modulation
$=T_{\mathrm{whirl}}(t_{o})+\frac{2\pi}{\bar{a}(t_{o})}\,\tilde{a}(t_{o})+\ldots,$ (IV.18)
where the second term arises from the logarithmic piece through $\ln s_{n+1}-\ln s_{n}=-2\pi/\bar{a}(t_{o})$. Linearizing Eq. (IV.18) using Eqs. (IV.7) and (IV.17) shows that $\Delta T_{n+1,n}$ acquires a harmonic modulation $\propto\mathrm{Re}\,{e^{-i\omega t_{o}}}$ with an amplitude set by $\gamma_{1},\tilde{a}_{1},a_{1}$.
Figure 6: Inter-image time delays $\Delta T_{n+1,n}(t_{o})$ from Eq. (IV.16). The baseline separation is set by the whirl period $T_{\rm whirl}=2\pi/\Omega_{\rm ph}$ with $\Omega_{\rm ph}=(3\sqrt{3}\,M)^{-1}$, while the weak modulation arises from the time-dependent coefficients in Eq. (IV.17) and the $\bar{a}(t_{o})$ dependence. The delays oscillate in phase with the spacing modulation $q(t_{o})$, reflecting their common origin in the near-photon-sphere kernel.
Figure 6 makes the phase-locked behavior explicit: variations in $\Delta T_{n+1,n}(t_{o})$ track the same QNM phase that controls $q(t_{o})$ and $\theta_{\rm ring}(t_{o})$. Eliminating $t_{o}$ between Eqs. (IV.18) and (IV.20) yields the differential correlation summarized in Eq. (V.12), further illustrating their common kernel origin.
Finally, the photon ring corresponds to the accumulation of the sequence $\theta_{n}(t_{o})$ as $n\to\infty$ [21]. From Eq. (IV.13) we obtain the instantaneous ring radius
@@
$=\theta_{c}\left[1+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\beta_{1}\right\}\right],$ (IV.19)
@@
$=e^{-2\pi}\left[1+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{2\pi\,a_{1}}{\bar{a}^{2}}\right\}\right]+\mathcal{O}(\varepsilon^{2}),$ (IV.20)
where we used $\bar{a}=1$ for Schwarzschild in the static limit. Thus, both the ring radius and the logarithmic spacing inherit the QNM frequency and phase from Eq. (III.7). The same analysis yields the magnifications $\mu_{n}(t_{o})$, which scale as $\mu_{n}\propto q(t_{o})^{n}$ up to a slowly varying prefactor; we refrain from writing the prefactor explicitly because it depends on the source configuration, while the time-dependent exponential factor follows directly from Eq. (IV.20). Related visibility-domain signatures and detectability of higher-order rings in realistic accretion models are discussed in [45]; an analytic gap parameter characterization of higher-order rings in spherically symmetric metrics is given in [7].
@@
Eqs. (IV.13)-(IV.20), together with the deflection law in Eq. (IV.6), provide a self-contained, analytic description of how a single even-parity $\ell=2,m=0$ QNM coherently imprints on the arrival times, multiplicity, and morphology of strong-field images, culminating in a phase-locked modulation of the photon ring.
@@ V Imaging Diagnostics and Continuity Across Regimes
We now assemble a single, observer-based description of lensing observables that is valid from the weak field to the time-dependent strong-deflection limit. The central object is the Born line-of-sight map in Eq. (III.7), which yielded the far-field expansion of Eq. (III.13) and the strong-field law of Eq. (IV.6). Our task is twofold: first, to show that these limits admit a common overlap where they agree term by term once expressed on the screen defined in Sec. II.2; second, to translate the matched deflection into gauge-safe imaging diagnostics for centroid motion, multi-image structure, and photon-ring modulations.
@@ V.1 The unified matching
The matching proceeds directly at the level of the master integral in Eq. (III.7). We split the integral at an arbitrary radius $r_{m}$ chosen so that $3M<r_{m}\ll r_{o}$, writing
@@
$+\int_{\lambda(r_{m})}^{\lambda_{o}}d\lambda\,\tilde{\mathcal{K}}_{A}\,e^{i\omega\tau}\equiv\mathcal{A}_{A}^{\text{in}}+\mathcal{A}_{A}^{\text{out}},$ (V.1)
@@
For $b\to b_{c}^{+}$ at fixed $r_{m}$, the inner integral reproduces the logarithmic structure derived in Eq. (IV.5), hence
@@
$+\mathcal{D}_{A}(\omega;r_{m})\Bigr]\biggr\}+\ldots,$ (V.2)
with $\mathcal{C}_{A}$ independent of $r_{m}$ and $\mathcal{D}_{A}$ carrying the $r_{m}$-dependence required to cancel that of the outer piece. Conversely, for $M/b\ll 1$ at fixed $r_{m}$, the outer contribution matches the far-field form in Eq. (III.10),
$\delta\alpha_{A}^{\text{out}}(b,t_{o})=\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{c_{A}(\omega;r_{m})}{b}\,\mathcal{F}_{A}(\omega b)\right\}+\mathcal{O}\,\left(\frac{M}{b^{2}}\right),$ (V.3)
where $c_{A}(\omega;r_{m})$ depends on the cutoff but $\mathcal{F}_{A}$ does not. Adding Eqs. (V.2)-(V.3) yields a representation that is uniformly valid for $b>b_{c}$.
Figure 8: Kernel split and $r_{m}$ independence of the sum. The master line-of-sight amplitude is decomposed as in Eq. (V.1) into the inner (near-photon-sphere) piece and the outer (far-zone) piece. Each contribution depends on the arbitrary split radius $r_{m}$ and carries complementary logarithmic/finite parts, yet their sum is independent of $r_{m}$. This illustrates that physical observables derived from the Born kernel are unique and do not depend on the cutoff.
Figure 8 shows how the $r_{m}$-dependence cancels between the inner and outer integrals in Eq. (V.1). This cancellation underlies the robustness of the matched coefficients in Eq. (V.4) and, ultimately, the overlap relation in Eq. (V.5).
To make contact with the time-dependent SDL coefficients of Eq. (IV.7), we expand the outer factor $\mathcal{F}_{A}(\omega b)$ for $b$ close to $b_{c}$. Using the standard near-critical relation between the longitudinal coordinate and the azimuthal angle, Eq. (IV.4), one finds that $\mathcal{F}_{A}$ carries a logarithmic dependence identical to that of the inner piece. The $r_{m}$-dependence cancels between $\mathcal{D}_{A}(\omega;r_{m})$ and $c_{A}(\omega;r_{m})$, and the sum can be cast as
@@
$+\frac{\beta_{1,A}(\omega)}{\frac{b}{b_{c}}-1}+b_{1,A}(\omega)\Bigr]\biggr\}+\ldots,$ (V.4)
which is the vectorial version of Eq. (IV.9) and fixes, component by component, the response coefficients $a_{1,A},\beta_{1,A},b_{1,A}$ in terms of the same kernel $\tilde{\mathcal{K}}_{A}$ used in the far-field calculation.
A practical matching condition follows by selecting an overlap domain where both asymptotics are valid. Let $b=b_{c}(1+\Delta)$ with $0<\Delta\ll 1$ and keep $M/b\ll 1$. Expanding Eq. (III.13) to first order in $\Delta$ and equating to Eq. (V.4) yields
@@
$+\bm{\Xi}(\omega),$ (V.5)
where $\bm{\mathcal{A}}=(c_{X}\mathcal{F}_{X},c_{Y}\mathcal{F}_{Y})$ from Eq. (III.13), $\hat{\bm{e}}_{b}$ is the screen radial unit vector, and $\bm{\Xi}(\omega)$ collects finite terms that are independent of the arbitrary split radius $r_{m}$. Expanding $A(b,\omega)$ about $b=b_{c}(1+\Delta)$ generates the $\ln\Delta$ and $1/\Delta$ structures that match the strong-deflection expansion on the inner side, making the regulator independence manifest. Eq. (V.5) states that the $1/b$ Born amplitude evaluated at $b_{c}$ encodes, through its singular $\Delta$-expansion, the three time-dependent SDL responses in Eq. (IV.7). In particular, the residue of the $1/\Delta$ pole fixes $\beta_{1}(\omega)$ and therefore the photon-ring modulation in Eq. (IV.19), while the coefficient of $\ln\Delta$ fixes $a_{1}(\omega)$ and hence the spacing modulation in Eq. (IV.20).
Figure 9: Matching weak and strong limits in the overlap domain. The weak-field Born expression (Eq. (III.13), dashed) and the linearized SDL law (Eq. (IV.9), solid) are compared at $b=b_{c}(1+\Delta)$ for small $\Delta>0$. Around an intermediate reference $\Delta$ the two curves coincide, and they agree within a band across the overlap window, illustrating Eq. (V.5): the near-critical expansion of the $1/b$ Born amplitude encodes the SDL coefficients, in particular the $1/\Delta$ residue that fixes $\beta_{1}(\omega)$.
Figure 9 visualizes the content of Eq. (V.5): the same kernel controls both limits, and their expansions meet smoothly for small but finite $\Delta$. The agreement pins down the triplet $(\beta_{1},a_{1},b_{1})$ that governs ring-radius and spacing modulations (Sec. (IV)).
Finally, since Eqs. (III.7), (III.13), and (IV.6) are all constructed using the observer tetrad of Eq. (II.9), the matching preserves the gauge safety guaranteed by Eq. (II.16). The result is a single kernel-based description in which weak-field centroid wobble, intermediate-field image shifts, and strong-field ring modulations are different faces of the same frequency-resolved response.
@@ V.2 Screen-plane observables: centroid, multi-images, and ring
All observables are defined on the screen introduced in Eq. (II.13) and inherit their time dependence from the master kernel in Eq. (III.7). We collect here the gauge-safe quantities that diagnose the weak, intermediate, and strong lensing regimes and emphasize how they are continuously related by the matching established in Sec. V.1.
@@
$\zeta\equiv\frac{X+iY}{r_{o}},$ (V.6)
so that a ray of impact parameter $b$ arriving at $(X,Y)$ corresponds to $\zeta(b,t_{o})=\alpha_{X}(b,t_{o})/r_{o}+i\alpha_{Y}(b,t_{o})/r_{o}$, with $\bm{\alpha}$ given by Eq. (III.3). For a narrow bundle around $b$ the centroid angle is the intensity-weighted mean,
$\zeta_{c}(t_{o})=\zeta_{c,0}+\varepsilon\,\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\mathcal{B}(\omega)\right\}+\mathcal{O}\,\left(\frac{M}{r_{o}}\frac{M}{b^{2}}\right),$ (V.7)
which is the complex form of Eq. (III.15). The complex amplitude $\mathcal{B}(\omega)$ depends on the bundle’s impact-parameter distribution through $\mathcal{A}_{A}(b,\omega)$ in Eq. (III.7). Since both the phase and modulus of $\mathcal{B}$ are determined by the same kernel that appears in the strong-field coefficients of Eq. (IV.7), the centroid wobble is phase-locked to the modulations of the relativistic images and of the photon ring discussed below. In particular, defining
$\Phi_{\rm c\,-\,ring}\equiv\arg\mathcal{B}-\arg\beta_{1},$ (V.8)
the relative phase between the weak-field centroid wobble and the strong-field ring-radius oscillation, Eq. (IV.19), is an overlap-domain diagnostic of the common QNM driver.
Figure 10: Phase-lock diagnostic across regimes. We show the phases associated with the weak-field centroid wobble (from Eq. (III.15)) and the strong-field ring-radius modulation (from Eq. (IV.19)), together with their difference $\Phi_{\rm c\!-\!ring}$ defined in Eq. (V.8). The two phases advance in parallel at frequency $\omega_{R}$ while their difference remains constant, demonstrating that weak- and strong-field screen observables are driven by the same QNM kernel.
Figure 10 exhibits the anticipated phase locking: the centroid and ring phases are parallel lines with a constant offset $\Phi_{\rm c\!-\!ring}$. This provides a direct, gauge-safe confirmation that the Born kernel controls both observables, as argued in Sec. V.1 and encoded in Eq. (V.5).
Relativistic images are labeled by the winding number $n\in\mathbb{N}$ and obey Eq. (IV.13). Their instantaneous angular radii $\theta_{n}(t_{o})$ separate naturally into a ring-centered offset,
@@
with $\theta_{\rm ring}(t_{o})$ given by Eq. (IV.19). The complex offset amplitude,
$\delta\theta_{n}^{\rm off}=\theta_{c}\,s_{n}^{(0)}\,\left(\frac{b_{1}}{\bar{a}}-\frac{\bar{b}-2\pi n}{\bar{a}^{2}}\,a_{1}\right),$ (V.10)
is the second term in Eq. (IV.14). The geometric decay $\propto s_{n}^{(0)}$ ensures that higher-order images accumulate at the ring while remaining coherently modulated at the QNM frequency $\omega$.
The logarithmic spacing is governed by $q(t_{o})$ in Eq. (IV.20). Writing $q(t_{o})=q_{0},[1+\varepsilon\,\mathrm{Re}{e^{-i\omega t_{o}}\,\delta_{q}}]$ with $q_{0}=e^{-2\pi/\bar{a}}$ and $\bar{a}=1$ for Schwarzschild, we have
$\delta_{q}=\frac{2\pi\,a_{1}}{\bar{a}^{2}},$ (V.11)
so that a measurement of the instantaneous ratio $\left(\Delta\theta_{n+1}/\Delta\theta_{n}\right)(t_{o})$ yields $\mathrm{Re}{e^{-i\omega t_{o}}\delta_{q}}$ directly. In combination with Eq. (IV.19), which determines $\mathrm{Re}{e^{-i\omega t_{o}}\beta_{1}}$, this provides two independent strong-field observables tied to the same kernel coefficients $a_{1}$ and $\beta_{1}$.
Arrival times follow from the travel-time function $T(b,t_{o})$ in Eq. (IV.15). Evaluated at $b_{n}(t_{o})$ from Eq. (IV.12), the inter-image delays $\Delta T_{n+1,n}(t_{o})$ are given by Eq. (IV.18) and thus oscillate at the QNM frequency through the modulations in Eq. (IV.17). Eliminating $t_{o}$ between Eqs. (IV.18) and (IV.20) gives a purely geometric correlation,
$\frac{d}{dt_{o}}\ln q(t_{o})=\frac{2\pi}{\bar{a}^{2}}\,\frac{d}{dt_{o}}\bar{a}(t_{o})=\mathcal{C}_{T}\,\frac{d}{dt_{o}}\Delta T_{n+1,n}(t_{o})+\mathcal{O}(\varepsilon^{2}),$ (V.12)
where $\mathcal{C}_{T}$ is a constant fixed by the static Schwarzschild values of $\tilde{a}$ and $\bar{a}$. Eq. (V.12) expresses the common origin of spacing and delay modulations in the near-photon-sphere kernel and, in our framework, is the time-domain analogue of the matching relation (V.5).
@@
$\mu_{n+1}(t_{o})\simeq q(t_{o})\,\mu_{n}(t_{o}),$ (V.13)
up to a slowly varying prefactor that is independent of $n$. Using Eq. (IV.20), the fractional modulation of $\mu_{n}$ at fixed $n$ is therefore $\delta\mu_{n}/\mu_{n}\simeq\mathrm{Re}{e^{-i\omega t_{o}}\delta_{q}}$, while at fixed $t_{o}$ the ratio $\mu_{n+1}/\mu_{n}$ isolates $q(t_{o})$. Together with Eqs. (IV.19)-(IV.20), this yields a closed set of strong-field observables $\theta_{\rm ring},q,\mu_{n+1}/\mu_{n}$} that are all phase-locked to $\omega$.
Finally, the weak- and strong-field diagnostics can be combined without reference to the arbitrary matching radius used in Eq. (V.1). A convenient pair is
$\left(\zeta_{c}(t_{o})-\zeta_{c,0}\right)\quad\text{and}\quad\theta_{\rm ring}(t_{o})-\theta_{c},$ (V.14)
whose complex phases are $\arg\mathcal{B}$ and $\arg\beta_{1}$, respectively. The constancy of $\Phi_{\rm c\,-\,ring}$ in Eq. (V.8) across the overlap domain provides a direct test of continuity from Eq. (III.13) to Eq. (IV.6). Because all quantities in Eqs. (V.7)-(V.14) are defined through the tetrad map Eq. (II.13), the gauge-independence argument of Eq. (II.16) applies verbatim: centroid motion, inter-image spacing, ring radius, and magnification ratios are bona fide $\mathcal{O}(\varepsilon)$ observables tied to a common first-order response encoded in the Born kernel (III.7).